Chapter 23: Homework Question 1: (1 point)

The following relates to Preston Company: Standard direct materials cost per unit: 2 pounds at $8.75 per pound. During the current year, Preston purchased and used 5,500 pounds of materials at a cost of $8 per pound to manufacture 2,500 units.

Required: Calculate the favorable amount of the materials price variance. When entering the answer in Blackboard, omit $ signs.

Chapter 23: Homework Question 2: (1 point) The following relates to Preston Company: Standard direct materials cost per unit: 2 pounds at $8.75 per pound. During the current year, Preston purchased and used 5,500 pounds of materials at a cost of $8 per pound to manufacture 2,500 units.

Required: Calculate the unfavorable amount of the materials quantity variance. When entering the answer in Blackboard, omit $ signs.

Chapter 23: Homework Question 3: (1 point) The Joy- Ride Corporation has established labor standards as 4 hours to manufacture one unit at a standard labor rate of $15 per hour. During the current year, 1,000 units were produced and 4,600 actual direct labor hours were recorded. Actual labor rate paid was $14.00 per hour.

Required: Calculate the favorable amount of the labor rate variance. When entering the answer in Blackboard, omit $ signs.

Chapter 23: Homework Question 4: (1 point) The Joy- Ride Corporation has established labor standards as 4 hours to manufacture one unit at a standard labor rate of $15 per hour. During the current year, 1,000 units were produced and 4,600 actual direct labor hours were recorded. Actual labor rate paid was $14.00 per hour.

Required: Calculate the unfavorable amount of the labor time variance.
Answer #1

1) Material price variance = (Standard price-actual price)actual quantity purchase = (8.75-8)*5500 = 4125 F

2) Material quantity variance = (Standard qty-actual qty)Standard price = (2500*2-5500)*8.75 = 4375 U

3) Labor rate variance = (Standard rate-actual rate)actual hour = (15-14)*4600 = 4600 F

4) Labor time variance = (Standard hour-actual hour)Standard rate = (4000-4600)*15 = 9000 U